Carla invested $28,000, part at 14% and part 9%. IF the total interest at the end of the year is $3,370, how much did she invest at 14%?
Answer by edjones(8007)

Let x be invested at 14% and 28000-x be invested at 9%.
.14x+.09(28000-x)=3370
.14x+2520-.09x=3370
.05x+2520-2520=3370-2520
.05x=850
.05x/.05=850/.05
x=$17,000
.
Check
17000*.14+11000*.09=3370
.
Ed